OpenStudy (anonymous):

rationalize denominator : please help 1.) 4/ sqrt. 7p 2.) srqt.98/x 3.) 9/3 sqrt.y 4.) sqrt.7 / sqrt.5 +7

OpenStudy (anonymous):

\[1.) 4/\sqrt{7p}\]

OpenStudy (anonymous):

2.) \[\sqrt{98}/x\]

OpenStudy (anonymous):

3.) \[9/ 3 \sqrt{y}\]

OpenStudy (anonymous):

4.) \[\sqrt{7} / \sqrt{5} + 7\]